# Quantum thermalization and the route to ergodicity

### v1

@@ Abstract
Contrast with Wigner theory.– Referring to the 2D energy landscape of the BHM, see Fig. 1, the width $\Delta x_{\rm shl}$ is finite, with negligible dependence on $\lambda$, see Fig. 2. This should be contrasted with the WBRM paradigm, where the horizonatal and the vertical dimensions are degenerate. Namely, $\Delta x_{\rm shl}=\Delta\varepsilon_{\rm shl}/\Delta_{0}$, where $\Delta_{0}$ is the mean level spacing. Thus, in the WBRM model, $x$ can be viewed as a dimensionless energy coordinate. This observation leads to the understanding that the mathematical-oriented WBRM model is an oversimplification. The actual route towards thermalization in a realistic physical system like the BHM is more complicated and non-universal.
@@
The route to ergodicity.– We identify five distinct physical regimes in the intricate progression of the BHM towards ergodicity, as the coupling strength $\lambda$ increases. These are marked by the shaded regions in Fig. 3, with a representative eigenstate for each in Fig. 4. For the general discussion we note that the relevant model parameters are the mean level spacing $\Delta_{0}$, the bandwidth $\Delta_{b}$, and the typical value $\bar{V}$ of the in-band matrix elements. It is useful to define dimensionless bandwidth ${b=\Delta_{E}/\Delta_{0}}$, and dimensionless coupling ${\tilde{\lambda}:=\lambda\bar{V}/\Delta_{0}}$.
@@
(a) The perturbative regime.– For very small values of $\lambda$, The peak probability $P_{\nu}$ is of order unity, and the $|E_{\nu}\rangle$ eigenstates can be described by 1st-order perturbation theory, which implies
@@
$p_{x,n}^{(\nu)}=\left|\frac{\left\langle x,n\middle|\lambda V\middle|x_{0},n_{0}\right\rangle}{\varepsilon_{x,n}-\varepsilon_{x_{0},n_{0}}}\right|^{2},\ \ \ (x,n)\neq(x_{0},n_{0})$ (13)
@@
(b) The Mixing regime.– As $\lambda$ is increased nearby levels are mixed, leading to level repulsion reflected by the spectral measure $r$. At this stage the eigenstates spread in the ’horizontal’ $x$ direction of the ${(x,\varepsilon)}$ energy shell, yet remain localized in the ’vertical’ $\varepsilon$ direction. A rough estimate for this horizontal mixing is provided by the standard Anderson model, predicting a localization length ${\xi\sim\lambda^{2}}$. Full horizontal mixing is attained once ${\xi\sim\sin N}$. This implies a border ${\tilde{\lambda}_{\rm mix}\sim\sqrt{N}}$ at which ’horizontal ergodicity’ $\overline{\Delta x_{\nu}}/\Delta x_{\rm shl}\approx 1$ is attained.
@@
(c) The Wigner regime.– Once the horizontal mixing is achieved, we get into the Wigner regime of vertical spreading, where the WBRM theory applies. A terse summary of the WBRM theory is provided in the [27]. The mixed levels form a core of vertical width ${\Gamma_{E}\propto\lambda^{2}}$. This core determines the participation number $\mathcal{M}_{\nu}$. Full ergodicity is reached once ${\Gamma_{E}}$ becomes comparable with the vertical width of the energy shell Eq. (11). Consequently, the expected ergodicity threshold is ${\tilde{\lambda}_{\rm erg}\sim\sqrt{b}}$.
@@
(d) The Ergodic regime.– For $\lambda>\lambda_{\rm erg}$ there is full mixing within the energy shell, both horizontally and vertically. This regime is semiclassical. To the extent it conforms to the WBRM model, we can say that in the Wigner regime ${\Gamma_{E}\ll\Delta_{E}<\Delta_{b}}$, while in the ergodic regime ${\Gamma_{E}\sim\Delta_{E}>\Delta_{b}}$, where $\Gamma_{E}$ indicates the width of the ‘core’ region of non-perturbative mixing.
@@
(e) The Integrable regime.– Finally, when ${\lambda\gg K}$ the intra-system hopping may be neglected and the system is reduced to the sum of disconnected integrable Bose-Hubbard dimers. Consequently, ergodicity is lost, $M_{\nu}$ and $\Delta\varepsilon_{\nu}$ decrease, and the state-to-state variance $\sigma_{x}$ increases as the ETH is violated. This regime is non-universal. In the WBRM the somewhat analogous regime is the Anderson regime ${\tilde{\lambda}>b^{3/2}}$, see [27] for details.
@@
False ergodicity measures.– Unlike $\overline{\mathcal{M}_{\nu}}/\mathcal{M}_{\rm shl}$, the ratio $\overline{\Delta\varepsilon_{\nu}}/\Delta\varepsilon_{\rm shl}$ can not be used as an ergodicity measure. In fact, the first equality in Eq. (11) implies that the latter ratio is of order unity, irrespective of ergodicity. A rather straightforward argument (see [27]) shows that ${\Delta\varepsilon_{\nu}}$ for any eigenstate is determined by the strength of the perturbation $V$, and is linear in $\lambda$. This is true even if the eigenstate is localized within the energy shell. The most dramatic demonstration is based on Eq. (13) for which ${\mathcal{M}\sim 1}$, but nevertheless the dispersion ${\Delta\varepsilon_{\nu}}$ is determined by the tails, and does not care about ${\mathcal{M}}$. Consequently, the linear $\lambda$ dependence is not affected by ${\tilde{\lambda}_{\rm mix}}$ neither by the ${\tilde{\lambda}_{\rm erg}}$ threshold, and cannot be used as an ergodicity measure.
@@
The regime borders.– Consider the scaling of the mixing- and ergodicity borders $\lambda_{\rm mix}$ and the $\lambda_{\rm erg}$, with $N$. Within a semiclassical treatment a rough estimate for the localization length is ${\xi=(1/\Delta_{0})D_{x}}$, where the diffusion coefficient in $x$ is ${D_{x}=N^{2}D_{cl}(\lambda)}$. Note that $D_{cl}\propto[\lambda V_{cl}]^{2}$, where $V_{cl}$ is the strength of perturbation in ‘classical’ units. The density of states is ${(1/\Delta_{0})=N^{d}\Omega_{cl}}$, where the number of degrees of freedoms is $d=M{-}1$, and $\Omega_{cl}$ is the phase-space area of the energy shell. We note that $\xi$ is possibly larger as discussed in [23]. From ${\xi<N}$ we get the condition
@@
Summary.– The thermalization of large disordered system can be modeled as the ergodization of small nearby chaotic subsystems, sometimes referred to as hot spots [26]. RMT modeling provides a partial understanding of the route towards ergodization. For physical subsytems a more careful definition of the energy-shell concept is required. Then it becomes clear that two distinct steps are involved, with different thresholds $\lambda_{\rm mix}$ and $\lambda_{\rm erg}$, that signify spectral quantum chaos and ETH ergodization respectively.

### v2

@@ VII Contrast with Wigner theory
Referring to the 2D energy landscape of the BHM, see Fig. 1, the width $\Delta x_{\rm shl}$ is finite, with negligible dependence on $\lambda$, see Fig. 2. This should be contrasted with the WBRM paradigm, where the horizonatal and the vertical dimensions are degenerate. Namely, $\Delta x_{\rm shl}=\Delta\varepsilon_{\rm shl}/\Delta_{0}$, where $\Delta_{0}$ is the mean level spacing. Thus, in the WBRM model, $x$ can be viewed as a dimensionless energy coordinate. This observation leads to the understanding that the mathematical-oriented WBRM model is an oversimplification. The actual route towards thermalization in a realistic physical system like the BHM is more complicated and non-universal.
VIII The route to ergodicity
@@
We identify five distinct physical regimes in the intricate progression of the BHM towards ergodicity, as the coupling strength $\lambda$ increases. These are marked by the shaded regions in Fig. 3, with a representative eigenstate for each in Fig. 4. For the general discussion we note that the relevant model parameters are the mean level spacing $\Delta_{0}$, the bandwidth $\Delta_{b}$, and the typical value $\bar{V}$ of the in-band matrix elements. It is useful to define dimensionless bandwidth $b=\Delta_{E}/\Delta_{0}$}, and dimensionless coupling $\tilde{\lambda}:=\lambda\bar{V}/\Delta_{0}$}.
@@
The perturbative regime.– For very small values of $\lambda$, The peak probability $P_{\nu}$ is of order unity, and the $|E_{\nu}\rangle$ eigenstates can be described by 1st-order perturbation theory, which implies
@@
$p_{x,n}^{(\nu)}=\left|\frac{\left\langle x,n\middle|\lambda V\middle|x_{0},n_{0}\right\rangle}{\varepsilon_{x,n}-\varepsilon_{x_{0},n_{0}}}\right|^{2},\ \ \ (x,n)\neq(x_{0},n_{0})\ \ \ \$(20)
@@
The Mixing regime.– As $\lambda$ is increased nearby levels are mixed, leading to level repulsion reflected by the spectral measure $r$. At this stage the eigenstates spread in the ’horizontal’ $x$ direction of the $(x,\varepsilon)$} energy shell, yet remain localized in the ’vertical’ $\varepsilon$ direction. A rough estimate for this horizontal mixing is provided by the standard Anderson model, predicting a localization length $\xi\sim\lambda^{2}$}. Full horizontal mixing is attained once $\xi\sim N$}. This implies a border $\tilde{\lambda}_{\rm mix}\sim\sqrt{N}$} at which ’horizontal ergodicity’ $\overline{\Delta x_{\nu}}/\Delta x_{\rm shl}\approx 1$ is attained.
@@
The Wigner regime.– Once the horizontal mixing is achieved, we get into the Wigner regime of vertical spreading, where the WBRM theory applies. A terse summary of the WBRM theory is provided in Appendix A. The mixed levels form a core of vertical width $\Gamma_{E}\propto\lambda^{2}$}. This core determines the participation number $\mathcal{M}_{\nu}$. Full ergodicity is reached once $\Gamma_{E}$} becomes comparable with the vertical width of the energy shell Eq. (11). Consequently, the expected ergodicity threshold is $\tilde{\lambda}_{\rm erg}\sim\sqrt{b}$}.
@@
The Ergodic regime.– For $\lambda>\lambda_{\rm erg}$ there is full mixing within the energy shell, both horizontally and vertically. This regime is semiclassical. To the extent it conforms to the WBRM model, we can say that in the Wigner regime $\Gamma_{E}\ll\Delta_{E}<\Delta_{b}$}, while in the ergodic regime $\Gamma_{E}\sim\Delta_{E}>\Delta_{b}$}, where $\Gamma_{E}$ indicates the width of the ‘core’ region of non-perturbative mixing.
@@
The Integrable regime.– Finally, when $\lambda\gg K$} the intra-system hopping may be neglected and the system is reduced to the sum of disconnected integrable Bose-Hubbard dimers. Consequently, ergodicity is lost, $M_{\nu}$ and $\Delta\varepsilon_{\nu}$ decrease, and the state-to-state variance $\sigma_{x}$ increases as the ETH is violated. This regime is non-universal. In the WBRM the somewhat analogous regime is the Anderson regime $\tilde{\lambda}>b^{3/2}$}, see Appendix A for details.
IX False ergodicity measures
@@
Unlike $\overline{\mathcal{M}_{\nu}}/\mathcal{M}_{\rm shl}$, the ratio $\overline{\Delta\varepsilon_{\nu}}/\Delta\varepsilon_{\rm shl}$ can not be used as an ergodicity measure. In fact, the first equality in Eq. (11) implies that the latter ratio is of order unity, irrespective of ergodicity. As discussed in Sec. (IV.2), a straightforward argument shows that $\Delta\varepsilon_{\nu}$} for any eigenstate is determined by the strength of the perturbation $V$, and is linear in $\lambda$. This is true even if the eigenstate is localized within the energy shell. The most dramatic demonstration is based on Eq. (20) for which $\mathcal{M}\sim 1$}, but nevertheless the dispersion $\Delta\varepsilon_{\nu}$} is determined by the tails, and is not affected by $\mathcal{M}$}. Consequently, the linear $\lambda$ dependence of $\overline{\Delta\varepsilon_{\nu}}$ is insensitive to both the $\tilde{\lambda}_{\rm mix}$} and the $\tilde{\lambda}_{\rm erg}$} thresholds, hence it cannot be used as an ergodicity measure.
X The regime borders
@@
Consider the scaling of the mixing- and ergodicity borders $\lambda_{\rm mix}$ and the $\lambda_{\rm erg}$, with $N$. Within a semiclassical treatment a rough estimate for the localization length is $\xi=(1/\Delta_{0})D_{x}$}, where the diffusion coefficient in $x$ is $D_{x}=N^{2}D_{cl}(\lambda)$}. Note that $D_{cl}\propto[\lambda V_{cl}]^{2}$, where $V_{cl}$ is the strength of perturbation in ‘classical’ units. The density of states is $(1/\Delta_{0})=N^{d}\Omega_{cl}$}, where the number of degrees of freedoms is d=M1d=M{-}1, and $\Omega_{cl}$ is the phase-space area of the energy shell. We note that $\xi$ is possibly larger as discussed in [30]. From $\xi<N$} we get the condition
@@
XI Summary
There are numerous studies where thermalization of one-dimensional spin-chains is considered [17, 18, 19, 20]. The perturbation $\lambda V$ in such models is responsible for non-integrability, and the length $L$ comes into consideration within a framework of a finite-size scaling scheme. The paradigm that we use for thermalization is quite different. It is motivated by the realization that thermalization of large disordered system can be modeled as the ergodization of small nearby chaotic subsystems, that exchange particles and energy, sometimes referred to as hot spots [37].
No special assumptions about the subsystem geometry are made in our model, and $L$ is not a physically meaningful geometrical parameter for the analysis. The number of sites $M_{\text{subsystem}}$ should be larger than 2 in order to have chaos, or better larger than 3 in order to avoid Kolmogorov-Arnold-Moser phase-space barriers (we used $M_{A}=M_{B}=4$}). The spectral structure illustrated in Fig. 1(b) as well as the distinction between horizontal and vertical ergodization demonstrated in Fig. 4 are not system-specific and apply to all bipartite composite systems, regardless of their geometry or size. This point is further demonstrated in Appendix B. What really counts is the Hilbert space dimension, and hence the number of particles ($N$) has appeared in our discussion.
RMT modeling provides a partial understanding of the route towards ergodization. For physical subsytems a more careful definition of the energy-shell concept is required. Then it becomes clear that two distinct steps are involved, with different thresholds $\lambda_{\rm mix}$ and $\lambda_{\rm erg}$, that signify spectral quantum chaos and ETH ergodization respectively.
@@ XI Summary
Appendix A The WBRM model
The WBRM can be written schematically as follows:
$H=\text{diag}\{\varepsilon_{n}\}+\text{offdiag}\{W_{n,m}\}$ (23)
Parameters:
$\Delta_{0}$ $=$ level spacing (24)
$\Delta_{b}$ $=$ range of hopping (25)
$\bar{W}$ $=$ strength of hopping (26)
Dimensionless parameters:
$b$ $=$ $\frac{\Delta_{b}}{\Delta_{0}}\ =\ \text{dimensionless band width}$ (27)
$\lambda$ $=$ $\frac{\bar{W}}{\Delta_{0}}\ =\ \text{dimensionless coupling}$ (28)
The Wigner regime is
$1<\lambda<\sqrt{b}$ (29)
In this regime there are two “width” scales. One is $\Gamma_{E}$ defined via PN calculation, and the other is $\Delta_{E}$ defined via a second moment calculation. One obtains
$\Gamma_{E}$ $=$ $2\pi\frac{\bar{W}^{2}}{\Delta_{0}}$ (30)
$\Delta_{E}$ $=$ $\sqrt{b}\bar{W}$ (31)
Accordingly,
$\frac{\Gamma_{E}}{\Delta_{0}}$ $=$ $2\pi\lambda^{2}\ =\ \text{dimensionless core width}$ (32)
$\frac{\Delta_{E}}{\Delta_{0}}$ $=$ $\sqrt{b}\lambda\ =\ \text{dimensionless shell width}$ (33)
In the non-ergodic Wigner regime $\Gamma_{E}\ll\Delta_{E}<\Delta_{b}$}, while in the ergodic regime $\Gamma_{E}\sim\Delta_{E}>\Delta_{b}$}. Note that in the RMT model there is also an Anderson localization regime $\lambda>b^{3/2}$}, where the $\Gamma_{E}/\Delta_{0}$ width saturates to the value $\Delta_{\xi}/\Delta_{0}=\xi\sim b^{2}$}, that is much smaller than the unbounded semiclassically expected result $\Delta_{E}/\Delta_{0}\propto\lambda$}.
Figure 5: Chaos and Ergodicity Measures. Same as Fig. 3 but with random internal couplings between all the sites of the coupled subsystems, selected from a normal distribution with zero mean and unit standard deviation. Thus, each subsystem is not a chain, but a fully connected cluster of arbitrary geometry.
Appendix B Irrelevance of geometry
In order to demonstrate the irrelevance of geometry in the context of our analysis, we have generated numerical results for coupled subsystems, where each subsystem is a fully connected cluster rather than a chain. The results are presented in Fig. 5. All the main features of Fig. 3 and, in particular, all five coupling regimes remain unchanged.